The degree measure of one of two supplementary angles is 6 more than one half that of the other. What are the degree measures of the angles?
Answer by Grinnell(63)

The key here is setting the problem up correctly!
The two angles add up to 180 degrees, Def. of Supp. Angles.
6+1/2x (6 more than 1/2 of the other)
x (the 'other' angle, OK?
6+1/2x+x=180--you with me?
3/2x=174
3x=348
x=116
6+1/2x=64